# largesymbols"0D largesymbols"0D

### v3

@@ 4.2 Monte Carlo studies
Truth Pretrained NNE SMLE
SMLE $R=50$
SMLE
SMLE
$R=50$ $\boldsymbol{\lambda}=(4,2)$
$\alpha_{0}$5.0 -5.009 (0.006) -5.779 (0.010) -4.696 (0.023) -3.347 (0.011) -2.767 (0.011)
@@
$\beta_{1}$0.6 -0.603 (0.002) -0.579 (0.003) -0.375 (0.005) -0.457 (0.005) -0.366 (0.005)
$\beta_{2}$0.5 -0.488 (0.002) -0.485 (0.003) -0.300 (0.005) -0.370 (0.005) -0.295 (0.005)
$\beta_{3}$0.3 -0.292 (0.002) -0.290 (0.003) -0.188 (0.005) -0.228 (0.006) -0.183 (0.005)
$\beta_{4}$1.0 -1.010 (0.003) -0.968 (0.003) -0.627 (0.005) -0.771 (0.005) -0.616 (0.005)
$\beta_{5}$0.4 -0.399 (0.003) -0.390 (0.003) -0.255 (0.005) -0.315 (0.006) -0.250 (0.005)
$\beta_{6}$0.5 -0.489 (0.003) -0.479 (0.003) -0.316 (0.006) -0.392 (0.006) -0.312 (0.005)
@@
Our first set of tests are Monte Carlo studies. Specifically, for each of the 12 real datasets, we: (i) keep the product, consumer, and advertising attributes, but (ii) replace the search and purchase choices with those simulated by the search model under a “true” $\boldsymbol{\theta}$. Then, we try to recover $\boldsymbol{\theta}$ from this dataset. Knowing the ground truth of $\boldsymbol{\theta}$ allows us to measure estimation accuracy. The true $\boldsymbol{\theta}$ is set to be close to its estimate later in Section 4.3.
@@
We first closely examine the results from the eCommerce: laptop dataset, reported in Table 4.2. Column 1 shows the true value of $\boldsymbol{\theta}$. Column 2 shows the estimate by the pretrained NNE. Columns 3 - 6 show the estimates by SMLE. The $R$ denotes the number of simulations per consumer in SMLE and $R=50$ is typical in literature. For the smoothed SMLE, $\boldsymbol{\lambda}$ collects the smoothing factors for search choices and for purchase choices. [^10] Some applications use the same smoothing factor for search and purchase choices. But doing so here worsens RMSE substantially. Whether multiple smoothing factors are necessary is generally application-dependent. The table shows three settings of $\boldsymbol{\lambda}$ to illustrate the sensitivity of SMLE’s estimate to $\boldsymbol{\lambda}$. We actually have conducted an expensive grid search of $\boldsymbol{\lambda}$, and $\boldsymbol{\lambda}=(4,1)$ gives the smallest root mean square error (RMSE) in this application.
@@
The upper plot of Figure 2 reports the RMSEs of the different estimators. The pretrained NNE has the lowest RMSE, outperforming even the SMLE that uses GHK and a very large $R=1000$. Although we generally expect MLE to be asymptotically efficient, the performance of SMLE is affected by simulation errors. In search models, simulating the likelihood is particularly difficult because of the large number of possible search and purchase combinations. Our result here suggests that this likelihood simulation remains challenging despite workarounds such as smoothing and GHK. The result also shows that pretrained NNE can achieve high estimation accuracy despite being off-the-shelf.
@@
The lower plot of Figure 2 reports the computation time needed for a single estimation run. All the times are measured on the same hardware. Recall that this dataset is a subsample from a single product category of the original multi-category data (see Section 4.1). [^11] In the original data, there are 624 categories. The laptop category has 41,918 search sessions. Our dataset is a subsample of 10,000 search sessions. Even so, the SMLE requires between 1.5 and 10 hours. The pretrained NNE takes about 1 second (so the orange bar is barely visible). This represents a time saving of 3 to 4 orders of magnitude. In addition, note that implementing SMLE requires additional costs not reflected in Figure 2, such as tuning smoothing factors, tuning optimization routine, trying starting points. Each of these tasks requires multiple estimation runs. Moreover, Figure 2 does not reflect the coding cost of SMLE, which can be substantial with GHK.
@@
Next, we expand the examination to all 12 datasets. The results are reported in Table 4. The table omits the smoothed SMLE because, as we have seen, its estimates are sensitive to the smoothing factors $\boldsymbol{\lambda}$ (and the optimal $\boldsymbol{\lambda}$ is application-dependent). The main finding with Table 4 is qualitatively the same as what we have seen with the eCommerce: laptop dataset. The pretrained NNE achieves either comparable or better estimation accuracy at drastically lower computational costs.
@@
It is worth noting that the sizes of the 12 datasets here are not large (see Table 2). In fact, recall that each dataset focuses on only a subset (a single product category, travel destination, or sales channel) of its source. Further, four of the 12 datasets have been subsampled to reduce the computation burden. Because computation time scales with data size, the absolute time savings on large datasets will be much greater than those indicated by Table 4.
@@ 4.3 Real-data estimates
Our second set of tests estimates the search model entirely from real data – that is, both real attributes and real consumer choices. We no longer have a ground truth for $\boldsymbol{\theta}$ and cannot directly measure estimation accuracy. Therefore, we assess whether the pretrained NNE’s estimates are sensible, by comparing them with the estimates from the GHK-based SMLE that uses the very large $R=1000$.
Figure 3 shows the estimates across the 12 datasets. Each plot corresponds to one dataset. The horizontal axis lists the individual parameters in $\boldsymbol{\theta}$. The blue bars are estimates by SMLE and the orange bars are estimates by pretrained NNE. A 95% confidence interval is shown on each bar. Overall, the estimates by pretrained NNE and by SMLE are closely matched.
@@ 5 Conclusion
This paper explores the feasibility of pretraining an off-the-shelf machine learning model to estimate a given structural econometric model. Importantly, we do not use the machine learning model to capture relationships in data. Instead, these relationships are prescribed by the structural model, and we use the machine learning model to recover this structural model’s parameter. Pretrained NNE enables users to estimate a structural model as easily as reduced-form regressions, while providing equal or better accuracy than existing estimation methods (at least in the context of the search model).
@@
Structural estimation is very valuable in empirical research, e.g., it offers insights into micro-founded consumer preferences and enables counterfactual analyses. However, potentially wider applications have been hindered by the high costs in both researcher effort and computation time. Pretrained estimators help remove these barriers.
@@
The ease of computation offered by pretrained NNE may facilitate new applications. One notable area is to integrate structural models with real-time algorithms, such as dynamic pricing, ranking algorithms, recommender systems, bandit experiments (e.g., Agrawal et al. 2019, Jain et al. 2024). Because structural models provide estimates of the latent underlying consumer preferences, these estimates can be fed into real-time algorithms to improve their decisions, e.g., what price to set, what items to recommend. Counterfactual analyses can be used to evaluate and provide feedback to the algorithms. However, structural estimation is time-consuming, which has made such applications difficult. Pretrained NNE offers a way to estimate structural models in real time, allowing more seamless integration of structural models into these algorithms.
@@
The intention of this paper is to explore the feasibility of pretrained estimators, rather than develop a comprehensive estimator for search models. This leaves areas where the pretrained NNE can be further improved. We note two potential areas. The first area is to increase the richness of the structural model. The search model considered in this paper is a standard yet simple one. Future work may allow for attributes that simultaneously affect utility and search costs, or allow for richer unobserved heterogeneity in utility and search costs. These extensions will require us to find additional reduced-form patterns to help identify the more complex structural model.
@@
A second area for improvement is to address potential endogeneity. For example, if researchers have an instrumental variable (IV) for price, how do we build a pretrained estimator that can leverage the instrument? At a high level, addressing this problem involves two tasks. The first task is to generate training datasets that include instruments. The second task is to pass relevant information regarding the instrument to the neural net. These tasks are not straightforward. However, incorporating IV in NNE will be a significant and valuable extension, because we frequently need to address endogeneity in structural estimation applications.
@@ Appendix A Data patterns
In Section 3.1 we have briefly described the data patterns $\boldsymbol{m}$ that feed into the neural net. Below, we provide the exact specification of $\boldsymbol{m}$.
We first recall some notation that will be needed to describe $\boldsymbol{m}$. Vectors $\boldsymbol{x}_{ij}^{(prod)}$, $\boldsymbol{x}_{ij}^{(ads)}$, and $\boldsymbol{x}_{i}^{(cons)}$ collect the product, advertising, and consumer attributes. It is assumed that all attributes are standardized. As before, let $d^{(prod)}$, $d^{(ads)}$, and $d^{(cons)}$ denote the numbers of attributes. These numbers are dataset-dependent. Let $d_{max}^{(prod)}$, $d_{max}^{(ads)}$, and $d_{max}^{(cons)}$ denote the maximum values that we currently set for the numbers of attributes (given in Table 1). These maximum values are not dataset-dependent.
We now define several consumer-level variables that will be used in $\boldsymbol{m}$. Recall that $J$ is the number of products per consumer and $n$ is the number of consumers.
@@
  • Let $\widetilde{y}_{i}^{(nf)}=\mathbb{I}\big{\{}\widetilde{y}_{i}^{(search)}>1\big{\}}$ indicates whether consumer $i$ did any non-free search.
  • Let $\widetilde{y}_{i}^{(buy)}=\sum_{j=1}^{J}y_{ij}^{(buy)}$, which is a dummy for whether consumer $i$ made a purchase.
@@
    The coefficients for $\{1,\;\boldsymbol{x}_{ij}^{(prod)},\boldsymbol{x}_{ij}^{(ads)},\boldsymbol{x}_{i}^{(cons)}\}$ in a logit regression of $y_{ij}^{(search)}$ on $\{1,\;\boldsymbol{x}_{ij}^{(prod)},\allowbreak\boldsymbol{x}_{ij}^{(ads)},\boldsymbol{x}_{i}^{(cons)},\bar{\boldsymbol{x}}_{i}^{(prod)},\bar{\boldsymbol{x}}_{i}^{(ads)}\}$. The non-intercept coefficients are padded with zeros and collected in a vector of fixed length $d_{max}$. The zero paddings enter the vector in the same way as in expression (4). This style of collecting regression coefficients applies to all the regressions below.
@@
All the coefficients above are estimated using ridge regression that imposes a small penalty on the coefficients. We find that using the ridge regression noticeably decreases the validation loss of the neural net (in predicting $\boldsymbol{\theta}$). We also find it helpful to use two separate penalty levels, and include the coefficients from both penalty levels in $\boldsymbol{m}$. Using two penalty levels further decreases the validation loss of the neural net. The two penalty levels we currently use are 0.001 and 1e-6. Using three or more penalty levels yields only marginal further decreases in the validation loss.
@@ Appendix B Detecting ill-suited applications
For an off-the-shelf tool that estimates a structural model, it is helpful for users to have a way to detect whether a dataset is ill-suited for the structural model. A real dataset is ill-suited if the underlying data generating process is poorly described described by the assumed structural model. If we apply the pretrained NNE to an ill-suited dataset, the real data patterns will likely be outside the range of data patterns used in pretraining. As a result, the neural net will be extrapolating and the estimates may not be meaningful.
Thus, we want to detect whether the $\boldsymbol{m}$ based on a given real dataset is atypical compared to the training examples $\{\boldsymbol{m}^{(\ell)}\}_{\ell=1}^{L}$. This task is not straightforward because $\boldsymbol{m}$ can have hundreds of elements, and these elements are not independent. Even if each individual element of $\boldsymbol{m}$ appears typical on its own, the vector $\boldsymbol{m}$ can still be atypical if it contradicts a relationship between multiple elements that holds in the training examples.
We propose a practical way to detect atypical $\boldsymbol{m}$, by making use of a tree-based model (boosted tree or random forest). In general, when a tree-based model and a neural net are trained on the same training examples, their predictions should not differ substantially within the region covered by training examples. However, the two models tend to extrapolate very differently. Specifically, let $\boldsymbol{f}^{NET}$ denote the neural net and $\boldsymbol{f}^{TREE}$ denote a tree-based model, both trained with the same training examples. We know $\boldsymbol{m}$ is atypical if $\boldsymbol{f}^{NET}(\boldsymbol{m})$ and $\boldsymbol{f}^{TREE}(\boldsymbol{m})$ differ substantially. In this case, we display a warning to users.
It is worth mentioning that $\boldsymbol{f}^{NET}$ has a lower validation loss than $\boldsymbol{f}^{TREE}$. This result suggests that, within the region covered by training examples, the neural net outperforms the tree-based model. Therefore, in applications that are not deemed ill-suited, we use the estimates from the neural net instead of the tree-based model.

### v4

@@ 4.2 Monte Carlo studies
Truth Pretrained NNE SMLE (GHK) $R=50$ SMLE (smoothed) $R=50$ $\boldsymbol{\lambda}=(4,1)$ SMLE (smoothed) $R=50$ $\boldsymbol{\lambda}=(2,1)$ SMLE (smoothed) $R=50$ $\boldsymbol{\lambda}=(4,2)$
$\alpha_{0}$ -5.0 -5.009 (0.006) -5.779 (0.010) -4.696 (0.023) -3.347 (0.011) -2.767 (0.011)
@@
$\beta_{1}$ -0.6 -0.603 (0.002) -0.579 (0.003) -0.375 (0.005) -0.457 (0.005) -0.366 (0.005)
$\beta_{2}$ -0.5 -0.488 (0.002) -0.485 (0.003) -0.300 (0.005) -0.370 (0.005) -0.295 (0.005)
$\beta_{3}$ -0.3 -0.292 (0.002) -0.290 (0.003) -0.188 (0.005) -0.228 (0.006) -0.183 (0.005)
$\beta_{4}$ -1.0 -1.010 (0.003) -0.968 (0.003) -0.627 (0.005) -0.771 (0.005) -0.616 (0.005)
$\beta_{5}$ -0.4 -0.399 (0.003) -0.390 (0.003) -0.255 (0.005) -0.315 (0.006) -0.250 (0.005)
$\beta_{6}$ -0.5 -0.489 (0.003) -0.479 (0.003) -0.316 (0.006) -0.392 (0.006) -0.312 (0.005)
@@
Our first set of tests are Monte Carlo studies. Specifically, for each of the 12 real datasets, we: (i) keep the product, consumer, and advertising attributes, but (ii) replace the search and purchase choices with the choices simulated by the search model under a “true” $\boldsymbol{\theta}$. Then, we try to recover $\boldsymbol{\theta}$ from this dataset. Knowing the ground truth of $\boldsymbol{\theta}$ allows us to measure estimation accuracy. We set the true $\boldsymbol{\theta}$ close to its estimate from Section 4.3.
@@
Below, we first closely examine the results from the eCommerce: laptop dataset. In Table 3, Column 1 shows the true value of $\boldsymbol{\theta}$, Column 2 shows the pretrained NNE’s estimate, and Columns 3 - 6 show the SMLE’s estimates. The $R$ denotes the number of simulations per consumer in SMLE and $R=50$ is typical in literature. For the smoothed SMLE, $\boldsymbol{\lambda}$ collects the smoothing factors for search choices and for purchase choices. [^10] Some applications use the same smoothing factor for search and purchase choices. But doing so here worsens RMSE substantially. Whether multiple smoothing factors are necessary is generally application-dependent. The table shows three settings of $\boldsymbol{\lambda}$ to illustrate the sensitivity of SMLE’s estimate to $\boldsymbol{\lambda}$. The optimal $\boldsymbol{\lambda}$ is generally application-dependent. We actually have conducted an expensive grid search of $\boldsymbol{\lambda}$, and $\boldsymbol{\lambda}=(4,1)$ gives the smallest root mean square error (RMSE) in this application.
@@
The upper plot of Figure 2 reports the RMSEs of the different estimators. The pretrained NNE has the lowest RMSE, outperforming even the GHK-based SMLE that uses a very large $R=1000$. Although we generally expect MLE to be asymptotically efficient, the performance of SMLE is affected by simulation errors. In search models, simulating the likelihood is particularly difficult because of the large number of possible search and purchase combinations. Our result here suggests that simulating this likelihood remains challenging despite workarounds such as smoothing and GHK. The result also shows that pretrained NNE can achieve high estimation accuracy despite being off-the-shelf.
@@
The lower plot of Figure 2 reports the computation time needed for a single estimation run. All the times are measured on the same hardware. Recall that this dataset is of moderate size (a subsample from a single product category; see Section 4.1). [^11] There are 624 categories in the original data. The laptop category has 41,918 search sessions. Our dataset is a subsample of 10,000 search sessions. Even so, the SMLE requires between 1.5 and 10 hours. The pretrained NNE takes about 1 second (so the orange bar is barely visible). This represents a time saving of 3 to 4 orders of magnitude. In addition, we note that implementing SMLE requires additional costs not reflected in Figure 2, such as tuning smoothing factors, tuning optimization routines, trying starting points. Each of these tasks requires multiple estimation runs. Moreover, the GHK takes considerable effort to code. While pretraining an NNE also has its own implementation costs, they arise only once during pretraining and are not borne by the subsequent users of the pretrained NNE.
@@
Next, we expand the examination to all 12 datasets. Table 4 reports the results. The table omits the smoothed SMLE because, as we have seen, its estimates are sensitive to the smoothing factors $\boldsymbol{\lambda}$ (and the optimal $\boldsymbol{\lambda}$ is application-dependent). The main finding with Table 4 is qualitatively the same as what we have seen with the eCommerce: laptop dataset. The pretrained NNE achieves either comparable or better estimation accuracy at drastically lower computational costs.
@@
Finally, it is worth noting that the 12 datasets here are not large (see Table 2). In fact, recall that each dataset focuses on only a subset (a single product category, travel destination, or sales channel) of its source, and four of the 12 datasets are subsampled to reduce the computation burden. Because computation time scales with data size, the absolute time savings on large datasets will be much greater than those indicated by Table 4.
@@ 4.3 Real-data estimates
Our second set of tests estimates the search model entirely from real data – that is, both real attributes and real consumer choices. We no longer have a ground truth for $\boldsymbol{\theta}$ and therefore cannot measure estimation accuracy directly. Instead, we assess whether the pretrained NNE’s estimates are sensible, by comparing them with the estimates from the GHK-based SMLE that uses the very large $R=1000$.
Figure 3 shows the estimates across the 12 datasets. Each plot corresponds to one dataset. The horizontal axis lists the individual parameters in $\boldsymbol{\theta}$. The orange bars show the pretrained NNE’s estimates, while the blue bars show the SMLE’s estimates. A 95% confidence interval is shown on each bar. Overall, we see that the estimates from the pretrained NNE and from the SMLE are closely matched.
@@ 5 Conclusion
This paper explores the feasibility of pretraining an off-the-shelf machine learning model to estimate a given structural econometric model. We do not use the machine learning model to capture relationships in data. Instead, these relationships are prescribed by the structural model, and we use the machine learning model to recover this structural model’s parameter. Pretrained NNE enables users to estimate a structural model as easily as reduced-form regressions, while providing equal or better accuracy than existing estimation methods (at least in the context of the search model).
@@
Structural estimation is very valuable in empirical research — it offers insights into micro-founded consumer preferences and enables counterfactual analyses. However, its wider use has been limited by the large amount of effort and time required to estimate structural models. Pretrained estimators remove these barriers, making structural models more accessible to both researchers and practitioners.
@@
Further, the ease of computation offered by pretrained NNE may facilitate new applications. One notable area is to integrate structural models with real-time algorithms, such as dynamic pricing, ranking algorithms, recommender systems, and bandit experiments (e.g., Agrawal et al. 2019, Jain et al. 2024). Because structural models provide estimates of latent consumer preferences, these estimates can be fed into real-time algorithms to improve their decisions, e.g., what price to set or what items to recommend. In addition, counterfactual analyses can be used to evaluate and provide feedback on the algorithms. However, in such applications the structural models must be estimated quickly in real time. Pretrained NNE offers a way to achieve such fast estimation.
@@
The intention of this paper is to explore the feasibility of pretrained estimators, rather than to develop a comprehensive estimator for search models. This leaves areas where the pretrained NNE can be further improved. We note two potential areas. The first area is to increase the richness of the structural model. The search model considered in this paper is a standard yet simple one. Future work may allow for attributes that simultaneously affect utility and search costs, or allow for richer unobserved heterogeneity in utility and search costs. These extensions will require us to find additional reduced-form patterns to help identify the more complex structural model.
@@
A second area for improvement is to address potential endogeneity. For example, if researchers have an instrumental variable (IV) for price, how do we build a pretrained estimator that can leverage the instrument? At a high level, addressing this problem requires two tasks. The first task is to generate training datasets that include instruments. The second task is to pass relevant information regarding the instrument to the neural net. These tasks are not straightforward. However, incorporating IV in NNE will be a significant and valuable extension, because we frequently need to address endogeneity in structural estimation applications.
@@ Appendix A Data patterns
Section 3.1 summarized the data patterns $\boldsymbol{m}$ that we feed into the neural net. Below, we provide the exact specification of $\boldsymbol{m}$.
We first recall some notation that will be needed. Vectors $\boldsymbol{x}_{ij}^{(prod)}$, $\boldsymbol{x}_{ij}^{(ads)}$, and $\boldsymbol{x}_{i}^{(cons)}$ collect the product, advertising, and consumer attributes, respectively. It is assumed that all attributes are standardized. As before, let $d^{(prod)}$, $d^{(ads)}$, and $d^{(cons)}$ denote the numbers of attributes. These numbers are dataset-dependent. Let $d_{max}^{(prod)}$, $d_{max}^{(ads)}$, and $d_{max}^{(cons)}$ denote the maximum values that we currently set for the numbers of attributes (given in Table 1). These maximum values are not dataset-dependent.
We now define several consumer-level variables that will be used. Recall that $J$ is the number of products per consumer and $n$ is the number of consumers.
@@
  • Let $\widetilde{y}_{i}^{(nf)}=\mathbb{I}\big\{\widetilde{y}_{i}^{(search)}>1\big\}$ indicate whether consumer $i$ did any non-free search.
  • Let $\widetilde{y}_{i}^{(buy)}=\sum_{j=1}^{J}y_{ij}^{(buy)}$, which is a dummy indicating whether consumer $i$ made a purchase.
@@
    The coefficients for $\{1,\;\boldsymbol{x}_{ij}^{(prod)},\boldsymbol{x}_{ij}^{(ads)},\boldsymbol{x}_{i}^{(cons)}\}$ in a logit regression of $y_{ij}^{(search)}$ on $\{1,\;\boldsymbol{x}_{ij}^{(prod)},\allowbreak\boldsymbol{x}_{ij}^{(ads)},\boldsymbol{x}_{i}^{(cons)},\bar{\boldsymbol{x}}_{i}^{(prod)},\bar{\boldsymbol{x}}_{i}^{(ads)}\}$. These coefficients are padded with zeros and collected in a vector of fixed length $d_{max}+1$. The paddings enter this vector in the same fashion as in expression (4). This style of collecting regression coefficients applies to all the regressions below.
@@
All the regressions above are estimated using ridge regression that imposes a small penalty on the coefficients. Using ridge regression leads to a noticeable decrease in the validation loss of the pretrained NNE. The validation loss is further decreased when we estimate each regression with two different penalty levels, and include the coefficients from both penalty levels in $\boldsymbol{m}$. We currently use the penalty levels of 0.001 and 1e-6. Using three or more penalty levels yields little additional decrease in the validation loss.
@@ Appendix B Detecting ill-suited applications
As discussed, we say a dataset is ill-suited if the associated data patterns $\boldsymbol{m}$ lies outside the range of data patterns seen in pretraining. In this case, the neural net has to extrapolate and its output may not be reliable. Because a pretrained NNE can be applied at low cost to many datasets, it is helpful to have a way to detect whether a particular dataset is ill-suited and to inform users when this is the case.
Specifically, we want to detect whether the $\boldsymbol{m}$ from a given real dataset is atypical (an outlier) relative to $\{\boldsymbol{m}^{(\ell)}\}_{\ell=1}^{L}$ from the training examples. We note that $\boldsymbol{m}$ has over a hundred dimensions, and the detection cannot be reduced to checking each dimension separately. Even if every individual element of $\boldsymbol{m}$ appears typical on its own, the vector $\boldsymbol{m}$ can still be atypical if it contradicts some relationship between elements that holds in the training examples.
We propose a practical method of detection by making use of a tree-based model (boosted tree or random forest). In general, when a tree-based model and a neural net are trained on the same training examples, their predictions should not differ substantially within the region covered by training examples. However, a tree-based model and a neural net tend to extrapolate very differently. So, let $\boldsymbol{f}^{NET}$ denote the neural net and $\boldsymbol{f}^{TREE}$ denote a tree-based model, both trained with the same training examples. We know $\boldsymbol{m}$ is atypical if $\boldsymbol{f}^{NET}(\boldsymbol{m})$ and $\boldsymbol{f}^{TREE}(\boldsymbol{m})$ differ substantially. In this case, we display a warning to users.
It is worth pointing out that $\boldsymbol{f}^{NET}$ has a lower validation loss than $\boldsymbol{f}^{TREE}$. In other words, within the region covered by training examples, the neural net outperforms the tree-based model. Therefore, in any application that is not deemed ill-suited, we use the output from the neural net instead of the tree-based model.
Appendix C Proof
Proposition 1 is proved by following Wei and Jiang (2024)’s proof (which builds on Chen 2007 and White 1989). The only difference is that Wei and Jiang (2024) work with a distribution of $(\boldsymbol{\theta},\boldsymbol{m})$ conditional on some observed $\boldsymbol{x}$, whereas Proposition 1 in the current paper considers the distribution of $(\boldsymbol{\theta},\boldsymbol{m})$ unconditional on $\boldsymbol{x}$. With this adjustment, the same argument in Wei and Jiang (2024) applies directly, so we omit repeating the details here.